A horizontal force of 200N is required to cause a 15kg block to slide up an  incline with an acceleration of . Find (a) the frictional force on the block and (b) the coefficient of friction



1
Expert's answer
2021-12-06T09:47:48-0500

Solution.

"F=200N;"

"m=15 kg;"

"\theta=20^o;"

"a=0.25 m\/s;"

"ma=Fcos\theta-mgsin \theta-F_f;"

"F_f=Fcos\theta-mgsin\theta-ma;"

"F_f=200\sdot0.94-15\sdot9.8\sdot0.34-15\sdot0.25=134.27N;"

"F_f=\mu N=\mu mgcos\theta\implies \mu=\dfrac{F_f}{mgcos\theta};"

"\mu=\dfrac{134.27}{15\sdot9.8\sdot0.94}=0.97;"


Answer:"a)F_f=134.27N;"

"b)\mu=0.97."